OpenStudy (anonymous):

Find the sum of the three expressions and choose the correct answer. -2u^3v + 5uv - 1 7u^3v - 2u^2v^2 5u^2v^2 - uv + 6 i got 5u^3v- 3u^2v^2+ 4uv- 5.

OpenStudy (anonymous):

-1 + 6 = ??

OpenStudy (anonymous):

Also: \[5u^2v^2 - 2u^2v^2 = ??\]

OpenStudy (anonymous):

Little mistakes of minus..

OpenStudy (anonymous):

-5u^3v- 3u^2v^2+ 4uv+ 7 ??

OpenStudy (anonymous):

No it is not correct..

OpenStudy (anonymous):

Your first and third term is right.. Second and fourth term have mistakes..

OpenStudy (anonymous):

waaa are yousure? wait..

OpenStudy (anonymous):

Take your time and do it again.. -1 + 6 = ??

OpenStudy (anonymous):

5

OpenStudy (anonymous):

Yes positive 5 and not -5..

OpenStudy (anonymous):

5u^2v^2 - 2u^2v^2 = ??

OpenStudy (anonymous):

5u^3v+ 3u^2v^2+ 4uv+ 5!!! sorry:)

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

Yes this is right..